OpenStudy (anonymous):

Ok this says to factor completely.. 12s^2+13s-35=???

OpenStudy (anonymous):

use the quadratic equation... -b+or- the square root of (b^2-4(a)(c)) all divided by 2(a)

OpenStudy (anonymous):

do you know this method? just wanna make sure you got it before i move on

OpenStudy (anonymous):

i have no clue lol

hero (hero):

12x^2+28x-15x-35 4x(3x+7)-5(3x+7) (3x+7)(4x-5)

hero (hero):

Says factor completely, not use quadratic formula

OpenStudy (anonymous):

use the polynomial given and label each section A, B, C. the 12 will be A the 13 will be B and the -35 will be C so now we have \[\left( \left( -13 \right)\pm \sqrt{\left( 13 \right)^2-4\left( 12 \right)\left( -35 \right)} \right)/2\left( 12 \right)\]